OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (anonymous):

please answer :(

OpenStudy (anonymous):

What is the slope of the perpendicular line to y = 2x + 2?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

The slope will have a negative reciprocal of the slope of y = 2x + 2

OpenStudy (anonymous):

the slope will be -1/2

OpenStudy (anonymous):

So you know that the equation will be y = -1/2x and the slope is -1/2 Now it has to pass through the point (6,3) Use the formula \(Y - Y_1 = -\frac{1}{2}(X - X_1) \) \(Y - 3 = -\frac{1}{2}(X - 6) \) <----can you solve

OpenStudy (anonymous):

yes okay so i add three on both sides

OpenStudy (anonymous):

no wait dont i distribute first?

OpenStudy (anonymous):

it would then be y-3=-1/2x -12 ??

OpenStudy (anonymous):

One sec I will write it out for you

OpenStudy (anonymous):

ohk

OpenStudy (anonymous):

 First you multiply each side by 2 to get ride of the fraction. the continue working the problem.